OpenStudy (anonymous):

Express the following equation in slope-intercept form: y + 41 = 11/4 (x + 16)

OpenStudy (anonymous):

Simplify equation first (distribute the 11/4). y + 41 = 11x/4 + 4 Subtract 41 from each side. y + 41 - 41 = 11x/4 + 4 - 41 y = 11x/4 - 37 Slope = 11/4 y-intercept = -37